OpenStudy (anonymous):

find the derivative of: arcsin(6x)

OpenStudy (amistre64):

im pretty sure its were you left it :)

OpenStudy (jamesj):

Let y = arcsin(6x). Then x = 1/6 . sin(y) Now, dx/dy = 1/6 . cos y = 1/6 sqrt(1 - sin^2 y) = 1/6 sqrt(1 - (6x)^2) Thus dy/dx = 6/sqrt( 1 - (6x)^2 )

OpenStudy (amistre64):

sin-1(6x) = y 6x =sin(y) 6 =cos(y) y' 6/cos(y) = y'